This is a differential equations course.

Find the general solution.

16. xt) = 1-30 2 1 -1 -2 -1 -1 0 o z(t)
Answer #1

x(H 1. o 11:0) 2 eigenvalues (Ardil - 0 O I 5.)?-412-74-6=0 . (*+2) (t?+28+3) = 0 ...2, = -1 + vai fgenuecta A-2 A- (-2) = ;Soluha y: center, ។ : (11.25 Solution for Getty Gerrit 9 - Vi e -Iri +- . et/icos eso con(0) [:-) -t Juu 2008(r2 t) + VI sin